OpenStudy (anonymous):

The line y = mx + m intersects the graph of y = x^2 at x = m + 2. Find the value of m.

OpenStudy (anonymous):

If you take the limit as x goes to infinity the top goes to x^2, and the bottom goes to x

OpenStudy (anonymous):

The function reduces to x^2/x=x, which has no vertical asymptote

OpenStudy (anonymous):

i think h.aysmp. y= -infinity

OpenStudy (anonymous):

Sorry, what I said was correct, but I meant horizontal, not vertical.

OpenStudy (anonymous):

x^2=mx+m

OpenStudy (anonymous):

(m+2)^2=m(m+2+1)

OpenStudy (anonymous):

m^2+4m+4=m^2+3m

OpenStudy (anonymous):

m=-4